OpenStudy (anonymous):

Use a proportion to solve the problem. Jim drove 250 miles in 5 hours. If he can keep the same pace, how long will it take him to drive 1,100 miles ? Help please

OpenStudy (anonymous):

\[\frac{250}{5}=\frac{1100}{?}\]

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

\[?=\frac{1100}{250} \times 5\]

OpenStudy (anonymous):

... that will make it easier...